A 4.55-kg ball of clay is thrown downward from a height of 3.51 m with a speed of 4.61 m/s onto a spring with k = 1730 N/m. The clay compresses the spring a certain maximum amount before momentarily stopping.

Find the maximum compression of the spring.

Find the total work done on the clay during the spring's compression
Answer #1

mass of the clay m=4.55 kg

speed v=4.61 m/sec

spring constant k=1730 N/m

height h=3.51 m

let compression is "x"

a)

by using energy equation,


m*g*(h+x)=1/2*k*x^2


4.55*9.8*(3.51+x)=1/2*(1730)*x^2


===> x=0.45m

compression x=0.45m =45 cm


b)


work done W=(1/2*k*x^2)

W=1/2*1730*0.45^2=175.16 J